OpenStudy (anonymous):

Help with negative exponents? 4. What is the value of 2x*-2y*-2 for x = 3 and y = -2? a. 1/18 b. 72 c. 2(-6)*-4 d. 1/648

OpenStudy (anonymous):

your question is \[2x ^{-2}*y ^{-2}\]

OpenStudy (anonymous):

yeah exactly. sorry, i wasn't sure how to write it out here.

OpenStudy (anonymous):

\[=\frac{ 2 }{ x ^{2}y ^{2} }=\frac{ 2 }{ \left( x*y \right)^{2} }=\frac{ 2 }{\left( 3*-2 \right)^{2} }=\frac{ 2 }{ 36 }=\frac{ 1 }{ 18 }\]

OpenStudy (anonymous):

thank you! it makes sense now.

OpenStudy (anonymous):

yw